OpenStudy (anonymous):

factor by grouping x^3-x^2-12x+12

OpenStudy (anonymous):

Solve for x over the real numbers: x^3-x^2-12 x+12 = 0 Factor the left hand side. The left hand side factors into a product with two terms: (x-1) (x^2-12) = 0 Solve each term in the product separately. Split into two equations: x-1 = 0 or x^2-12 = 0 Look at the first equation: Solve for x. Add 1 to both sides: x = 1 or x^2-12 = 0 Look at the second equation: Isolate terms with x to the left hand side. Add 12 to both sides: x = 1 or x^2 = 12 Eliminate the exponent on the left hand side. Take the square root of both sides: Answer: | | x = 1 or x = 2 sqrt(3) or x = -2 sqrt(3)